OpenStudy (anonymous):

(15q^6+5q^2)(5q^4)^-1

sam (.sam.):

(15q^(6)+5q^(2))(5q^(4))^(-1) 1/(5q^(4))*(15q^(6)+5q^(2)) (5q^(2)(3q^(4))+5q^(2)(1))/(5q^(4)) 5q^(2)(3q^(4)+1)/5q^(4) (3q^(4)+1)/q^(2)

OpenStudy (anonymous):

Get rid of negative exponent: (5q^4)^-1 = 1/ 5q^4 (15q^6+5q^2) / 5q^4 Factor out to simplify: = 5 q^2 ( 3 q^4 + 1) / 5q^4 = ( 3 q^4 + 1)/ q^2